OpenStudy (anonymous):

Tell whether the lines for each pair of equations are parallel, perpendicular, or neither y=-7/8x – 1 and 32x – 28y = –36

OpenStudy (karatechopper):

neither

OpenStudy (anonymous):

for perpendicular lines the slopes would multiply to give -1 and for parallel lines the slopes will be equal

OpenStudy (anonymous):

they are perpendicular y=-7/8x-1 32x-28y=-36 y=8/7x +9/7 mulitply the slopes -7/8 x 8/7 =-1

OpenStudy (akshay_budhkar):

yea as Alisha22 says we have to find both the slopes. in the first equation of the line we can see it is -7/8 comparing the equ. to y=mx+c in the second 1 we have the sleope by rearranging 28y=32x+36 therefore y=8/7x+36/28 so slope is 8/7 the product is -1 so they are perpendicular to each other